OpenStudy (anonymous):

Verify the identity. (7 points) cos(x+pi/2) = -sin x

OpenStudy (solomonzelman):

use the following rule: \(\large\color{slate}{\cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b) }\)

OpenStudy (anonymous):

@SolomonZelman so can you help me